# Trapped-ion Laser Cooling in Structured Light Fields

### v2

@@ Abstract
Laser cooling is fundamental to quantum computation and metrology with trapped ions, and can occupy a majority of runtime in current systems. A key limitation to cooling arises from unwanted carrier excitation, which in typically used running wave (RW) fields invariably accompanies the sideband transitions effecting cooling. We consider laser cooling in structured light profiles enabling selective sideband excitation with nulled carrier drive; motivated by integrated photonic approaches’ passive phase and amplitude stability, we propose simple configurations realizable with waveguide addressing using either standing wave (SW) or first-order Hermite-Gauss (HG) modes. We quantify performance of Doppler cooling from beyond the Lamb-Dicke regime (LDR), and ground-state (GS) cooling using electromagnetically induced transparency (EIT) leveraging these field profiles. Carrier-free EIT offers significant benefits simultaneously in cooling rate, motional frequency bandwidth, and final phonon number. Carrier-free Doppler cooling’s advantage is significantly compromised beyond the LDR but continues to hold, indicating such configurations are applicable for highly excited ions. Our simulations focus on level structure relevant to ${}^{40}\text{Ca}^{+}{}$, though the carrier-free configurations can be generally applied to other species. We also quantify performance limitations due to polarization and modal impurities relevant to experimental implementation. Our results indicate potential for simple structured light profiles to alleviate bottlenecks in laser cooling, and for scalable photonic devices to improve basic operation quality in trapped-ion systems.
@@
Trapped ions constitute a leading platform for quantum computation (QC) Bruzewicz et al. (2019) and simulation, with recent systems demonstrating algorithms on systems of tens of qubits with arbitrary connectivity Moses et al. (2023); Postler et al. (2024), leveraging the high-fidelity entangling gates Ballance et al. (2016); Gaebler et al. (2016); Clark et al. (2021); Löschnauer et al. (2024), state preparation, measurement, and readout Harty et al. (2014) achievable in this platform. The high coherence and precise control of systematics achievable in ion traps also enables timekeeping at the present limits of accuracy Brewer et al. (2019); Hausser et al. (2024).
@@
High-fidelity trapped-ion control in QC systems generally requires laser cooling to near the motional ground state (GS) Leibfried et al. (2003), also of interest for next-generation clocks Kulosa et al. (2023). To this end, Doppler cooling to near the Doppler limit may be followed by electromagnetically induced transparency (EIT) cooling Morigi et al. (2000a); Roos et al. (2000); Eschner et al. (2003), polarization gradient cooling (PGC) Birkl et al. (1994), and/or sideband cooling (SBC) Diedrich et al. (1989), to successively bring expectation phonon number $\bar{n}$ to near the GS Leibfried et al. (2003). Due to inevitable heating of atoms during operation De Leon et al. (2021) and because readout and transport operations repeated in trapped-ion QC operation Kielpinski et al. (2002) can excite ion motion, laser cooling generally must be repeatedly applied. Given the cooling rates achievable, laser cooling can occupy a majority of runtime in current trapped-ion QC systems Pino et al. (2021). Since operation times are a key limitation of trapped-ion systems, cooling schemes with improved cooling rate and final phonon number may address a crucial bottleneck in trapped-ion systems.
@@
Figure 1: (a) Steady state excitation population of a two-level system as a function of detuning $\Delta$. Carrier, BSB, and RSB transitions correspond to black, blue, and red dotted lines respectively. Due to power broadening, higher beam power ($s_{0}=2\Omega_{0}^{2}/\Gamma^{2}$) reduces the contrast between the RSB and BSB scattering rates, compromising cooling performance. (b) Probability distribution of a ${}^{40}\text{Ca}^{+}$ ion in thermal states for $\omega_{m}=2\pi\times 3$ MHz, centered at the nodal point of a $\lambda=397$ nm SW and a $w_{0}=500$ nm TE10 mode. The wave function of higher phonon number samples a larger field region, suggesting different cooling behaviors when outside/inside LDR. (c) Two experimental schemes for carrier-suppressed driving using integrated photonics, based on two counter-propagating beams from the same input to form a passively phase-stable SW, or a TE10 HG mode delivered from a single grating emitter fed by a TE1 guided mode Beck et al. (2024). The cooling beam’s intensity profiles are sketched on the right. (d) The eight-level structure relevant to ${}^{40}\text{Ca}^{+}$ ($S_{1/2},P_{1/2},D_{3/2}$) with lasers for Doppler cooling. $\Delta_{c},\Delta_{r}$ are detuning between $S_{1/2}$ and $P_{1/2}$, $D_{3/2}$ and $P_{1/2}$ respectively at zero magnetic field.
@@
Key limitations to cooling performance arise from couplings that accompany the desired interactions when employing the usual plane-wave-like running wave (RW) fields. In a frame rotating with laser frequency $\omega$, interaction between a RW laser field and an atom represented as a two-level system with ground and excited states $|g\rangle$ and $|e\rangle$, oscillating in a single motional mode at angular frequency $\omega_{m}$ can be described with a dipole interaction Hamiltonian expanded as:
@@
$\langle e|\hat{V}_{\mathrm{dip}}|g\rangle=\frac{\Omega_{0}}{2}e^{i\mathbf{k}\cdot\mathbf{\hat{R}}}=\frac{\Omega_{0}}{2}(1+i\eta(\hat{a}+\hat{a}^{\dagger})+...),$ (1)
@@
for wave vector $\mathbf{k}$ with $\left\lvert\mathbf{k}\right\rvert=k_{0}$ in free space, Rabi frequency $\Omega_{0}$, position operator $\mathbf{\hat{R}}=\mathbf{x_{0}}(\hat{a}+\hat{a}^{\dagger})$, and LD parameter $\eta=\eta_{0}\cos(\theta)$ where $\eta_{0}=k_{0}x_{0}=k_{0}\sqrt{\frac{\hbar}{2m\omega_{m}}}$ and $\theta$ is the angle between $\mathbf{k}$ and $\mathbf{x_{0}}$. The second term $\propto\eta$ describes coupling to motional sidebands due to first-order field gradients desired for cooling; larger scattering rate on the red sideband (RSB) as compared to the blue sideband (BSB) leads to cooling (Fig. 1). The first order term, due to the field at the ion location, represents carrier coupling that does not affect the motional state. Since typically $\eta<1$, the carrier drive strength exceeds the sideband and is problematic for two reasons. As $\Omega_{0}$ is increased to cool more rapidly, this carrier excitation saturates the internal population, reducing contrast in the excitation rate between the RSB and BSB, thereby limiting cooling rates and compromising the final phonon number achievable (Fig. 1a). Additionally, while carrier excitation is associated with no motional excitation, the subsequent spontaneous decay results in net heating. Thus, carrier excitation limits achievable cooling rate and additionally causes net heating.
@@
These limitations can be alleviated via suppression of carrier excitation and selective sideband coupling achievable at appropriate locations in structured light fields Cirac et al. (1992); Zhang et al. (2012); deLaubenfels et al. (2015); Vasquez et al. (2023). For electric dipole transitions within atoms trapped at points with zero intensity and nonzero field gradient, e.g. at SW nodes or nodal lines of higher-order Hermite-Gaussian (HG) modes (Fig. 1b), the carrier term (and all even orders of the expansion in Eq. 1) is nulled, with the first sideband term the leading contribution.
@@
Recent development of integrated waveguide photonics for optical control in planar ion traps Mehta et al. (2016); Niffenegger et al. (2020); Mehta et al. (2020); Vasquez et al. (2023); Ivory et al. (2021); Kwon et al. (2024); Mordini et al. (2024) offers a route to realize these mechanisms without the phase drifts and pointing instability that pose significant challenges in conventional free-space implementations Schmiegelow et al. (2016). For example, SW generation straightforwardly leverages counterpropagating beams derived from a single waveguide input (Fig. 1c, top) as demonstrated in Vasquez et al. (2023), and first-order HG mode generation can be achieved via emission of higher-order waveguide modes, e.g. a TE10 mode Beck et al. (2024) allowing use of transverse field gradients at the intensity node along the beam center (Fig. 1c, bottom). While the form of the interaction Hamiltonian is the same to lowest order for these two cases, the effective LD parameter in the latter case, $\eta_{\text{HG}}=\frac{2\sqrt{2}}{w_{0}}x_{0}$ West et al. (2021) reflects motional coupling from transverse field gradients, additionally offering avenues to engineer interactions to accommodate different ion crystal configurations and motional mode orientations. While we focus on these relatively simple field configurations, similar considerations would apply to more general structured light profiles Verde et al. (2023).
@@
In this work, we examine the practical performance of Doppler and EIT cooling using structured light fields through master equation simulation. We consider ${}^{40}\text{Ca}^{+}$’s relevant eight-level structure ($S_{1/2},P_{1/2},D_{3/2}$) in Fig. 1d, although the methods analyzed are applicable to other species. Our treatment extends previous theoretical work in considering Doppler cooling from well beyond the LDR (Fig. 1b), and accounting for sideband saturation effects.
@@
In the results presented below we show significant enhancement possible for carrier-free EIT GS cooling, simultaneously in cooling rate, motional mode frequency bandwidth, and over an order of magnitude reduction in final phonon number $\bar{n}_{ss}$ achievable. We quantify the sensitivity to these enhancements to polarization impurities, a likely experimental limitation. Realization of these simultaneous advantages may significantly reduce the SBC required to achieve GS occupancies required for high-fidelity control. For Doppler cooling, despite the expected degradation in SW cooling of highly excited ions as compared to within the LDR, we still find that a modest enhancement for SW cooling obtains. While sub-Doppler cooling via PGC also can leverage integrated photonic beam delivery Hattori et al. (2022) and benefit from ion positioning at polarization SW nodes, the expected enhancement in $\bar{n}_{ss}$ compared to with RWs is 2×2$\times$2 × Joshi et al. (2020). Given the more significant gains predicted for EIT here, we focus on EIT as a sub-Doppler and GS cooling method. We present our results for the simpler Doppler followed by EIT cooling below, commenting on the impact of $\eta_{0}$ on both schemes. This work indicates potential for scalable photonic hardware to alleviate bottlenecks in trapped-ion control, focusing here on laser cooling, and informs experimental implementation in integrated devices.
@@ I Carrier-free Doppler cooling
The treatment of Cirac et al. (1992) showed that Doppler cooling for ions within the LDR at SW nodes allows higher cooling rates via use of higher cooling beam intensities, as well as lower final $\bar{n}_{ss}$, than possible with RWs. Since SW cooling is limited by saturation on sideband transitions ($\Omega_{\mathrm{SB}}\propto\eta\Omega_{0}$) instead of the carrier ($\Omega_{\mathrm{car}}=\Omega_{0}$) as in the usual RW, in a SW configuration and for ions within the LDR, $W_{c}\propto\Omega_{\mathrm{SB}}^{2}$ can be increased by of order ${\sim 1}/\eta^{2}$ before saturating the two-level system. This delayed saturation enables a significant gain in $W_{c}$ given that typically, $\eta\sim 0.1$. Furthermore, an approximately 2×2$\times$2 × reduction in final phonon number $\bar{n}_{ss}$ is predicted.
@@
However, significantly different behavior, including heating, appears at other positions in the SW profile Cirac et al. (1992). Wavefunctions of highly excited ions outside the LDR, of practical interest to Doppler cooling, span lengthscales comparable to the wavelength (Fig. 1b), and hence compromised SW cooling performance is expected as compared to within the LDR as treated in Cirac et al. (1992). To study the cooling behavior beyond the LDR, here we employ numerical Lindblad master equation simulations including higher-order expansion terms (see Appendix). Our simulation quantifies the enhancement of Doppler cooling using SWs as compared to RWs both outside and inside the LDR, including effects of saturation from sideband excitation. We also discuss experimental configurations of carrier-free Doppler cooling based on integrated photonic addressing Vasquez et al. (2023); Mehta et al. (2020).
@@
In particular, we simulate the eight-level ${}^{40}\text{Ca}^{+}$ structure as an example with a $\hat{\pi}$-polarized cooling beam ($\Omega_{c}$, $\Delta_{c}$) addressing the $S_{1/2}\leftrightarrow P_{1/2}$ transition and a repumper ($\Omega_{r}=2\pi\times 10$ MHz, $\Delta_{r}$) with polarization $\hat{\epsilon}_{p}=\hat{\pi}/\sqrt{2}+(\hat{\sigma}_{+}-\hat{\sigma}_{-})/2$ between $D_{3/2}\leftrightarrow P_{1/2}$ as shown in Fig. 1a, and with a quantizing magnetic field of $B=1$ G. Time evolution of the expectation phonon number $\bar{n}(t)$ starting from a thermal state $\bar{n}(0)$ is extracted from master equation evolution. The cooling rate at a given $\bar{n}$ is defined as $W_{c}=-\dot{\bar{n}}/\bar{n}$, and the cooling limit $\bar{n}_{ss}$ the final steady-state expectation phonon number.
@@
Figure 2: (a) Cooling rate (line) and limit (point) for $\omega_{m}=2\pi\times 3$ MHz and $\Delta_{c}=-\Gamma/2$. EF$=1.5$ for $n=50$ (solid line), EF$=4.9$ for $n=5$ (dotted line). The additional D states decrease the cooling for both SW and RW but also reduce the RW carrier heating. (b) Time evolution of SW and RW cooling for $\omega_{m}=2\pi\times 3$ MHz. Using $\Delta_{c}=-0.5\Gamma$ and starting at $\bar{n}=50$, SW Doppler cooling reaches $\bar{n}=4$ within $34~{}\mu s$ with optimal Rabi frequency $\Omega_{sw}=1.8\Gamma$, while RW Doppler cooling takes $75~{}\mu s$ with optimal $\Omega_{rw}=0.8\Gamma$.
@@
Fig. 2a shows $W_{c}$ in the full eight-level system as a function of $\Omega_{c}$ for $\omega_{m}=2\pi\times 3$ MHz. $\Delta_{c}$ is set to $-\Gamma_{c}/2$ which optimizes the cooling limit $\bar{n}_{ss}$. For small $\Omega_{0}$, RW and SW Doppler cooling rates increase quadratically with $\Omega_{0}$ both inside ($\bar{n}=5$) and outside the LDR ($\bar{n}=50$) before saturation. For higher $\bar{n}$, the EF drops but remains larger than 1, achieved around the same $\Omega_{c}$ both outside and near the LDR. Combined with the consistent improvement in $\bar{n}_{ss}$, this validates the advantage of SW Doppler cooling in the experiments. In Fig. 2b, we simulate a complete Doppler cooling process from $\bar{n}=50$ outside the LDR to the Doppler limit, using the optimal parameters for cooling rate found in Fig. 2a. We observe SW Doppler cooling take less than half of the time as compared to RW to reach $\bar{n}=4$, a reasonable starting point for ground-state cooling.
Experimental implementation of carrier-free Doppler cooling may employ the configurations shown in Fig. 1c. Though for practical beam waists, the TE10’s effective $\eta$ is lower than in a RW, optimal sideband coupling can be recovered by increasing $\Omega_{0}$ inside the LDR to hold $\eta\Omega_{0}$ constant. This allows cooling radial modes of multi-ion strings without requiring precise positioning of multiple ions along a series of SW fringes. Outside the LDR, however, cooling through higher-order sideband excitations (with coupling strengths proportional to $\eta^{k}$ for a $k$th order sideband) contribute significantly to cooling, which are attenuated for lower driving beam LD parameter $\eta$ and fixed first-order coupling $\eta\Omega$. In contrast the heating through spontaneous decay on a $k$th order sideband is proportional to $\eta_{0}^{k}$ which is independent of the driving beam profile. As a result, the EF in cooling rate achievable with SW cooling decreases outside the LDR when using lower $\eta$ values. For example, in our eight-level system simulations we observe that the EF for SW cooling with $\eta=\eta_{0}/4$ compared to the best case RW with $\eta=\eta_{0}$ drops to nearly unity for $\bar{n}=50$.
Our results quantify the performance of carrier-free Doppler cooling beyond the LDR previously treated. Though using ${}^{40}\text{Ca}^{+}$ in the simulation as an example, similar advantage would hold for other ion species and are more favorable for lower intrinsic $\eta_{0}$ Xing and Mehta (2023). While the advantage offered is modest in comparison to well within the LDR, our results indicate that such configurations can still be applied for highly excited ions. Therefore, early cooling stages can share the same cooling beam profile with carrier-free GS cooling schemes, for which more substantial enhancements are predicted.
II Carrier-free GS EIT Cooling

### v3

@@ Abstract
Laser cooling is fundamental to quantum computation and metrology with trapped ions. In particular, quantum ground state cooling required for many gate implementations can occupy a majority of runtime in current systems. A key limitation to cooling arises from unwanted carrier excitation, which in typically used running wave (RW) fields invariably accompanies the sideband excitations effecting cooling. Laser cooling in structured light profiles enables selective sideband excitation with nulled carrier drive; motivated by integrated photonic approaches’ passive phase and amplitude stability, we propose simple configurations realizable with waveguide addressing using either standing wave (SW) or first-order Hermite-Gauss modes. We quantify performance of Doppler cooling from beyond the Lamb-Dicke (LD) regime, and ground-state cooling using electromagnetically induced transparency (EIT) leveraging these field profiles. Carrier-free EIT offers significant benefits in cooling rate and motional frequency bandwidth, and final phonon numbers comparable to those achievable with resolved sideband cooling. Carrier-free Doppler cooling’s advantage is significantly compromised beyond the LD regime but continues to hold, indicating such configurations are applicable for highly excited ions. Our simulations focus on level structure relevant to ${}^{40}\text{Ca}^{+}$, though the carrier-free configurations can be generally applied to other species. We also quantify performance limitations due to polarization and modal impurities relevant to experimental implementation. Our results indicate potential for simple structured light profiles to alleviate bottlenecks in laser cooling, and for scalable photonic devices to improve basic operation quality in trapped-ion systems.
@@
Trapped ions constitute a leading platform for quantum computation [1] and simulation, with recent systems demonstrating algorithms on tens of qubits with arbitrary connectivity [2, 3], leveraging the high-fidelity entangling gates [4, 5, 6, 7], state preparation, measurement, and readout [8] achievable in this platform. The high coherence and precise control of systematics achievable in ion traps also enables timekeeping at the present limits of accuracy [9, 10].
@@
High-fidelity trapped-ion control in quantum computing systems typically requires laser cooling to near the motional ground state [11], also of interest for next-generation clocks [12]. To this end, Doppler cooling may be followed by electromagnetically induced transparency (EIT) cooling [13, 14, 15], polarization gradient cooling [16], and/or sideband cooling [17, 18], to successively bring expectation phonon number $\bar{n}$ to near zero [11]. Due to inevitable heating of atoms during operation [19] and because readout and transport operations repeated in trapped-ion quantum computing operation [20] can excite ion motion, ground-state cooling generally must be repeatedly applied and can occupy a majority of runtime in current trapped-ion quantum computing systems [21]. Since operation times are a key limitation of trapped-ion systems, cooling schemes with improved cooling rate and final phonon number may address a crucial bottleneck in trapped-ion systems.
@@
Figure 1: (a) Steady state excitation population of a two-level system as a function of detuning $\Delta$. Carrier $|g,n\rangle\rightarrow|e,n\rangle$, blue sideband $|g,n\rangle\rightarrow|e,n+1\rangle$, and red sideband $|g,n\rangle\rightarrow|e,n-1\rangle$ excitations correspond to black, blue, and red dotted lines respectively. Due to power broadening, higher beam power ($s_{0}=2\Omega_{0}^{2}/\Gamma^{2}$) reduces the contrast between the red sideband and blue sideband scattering rates, compromising cooling performance. (b) Probability distribution of a ${}^{40}\text{Ca}^{+}$ ion in thermal states of different $\bar{n}$ for $\omega_{m}=2\pi\times 3$ MHz, centered at the nodal point of a $\lambda=397$ nm SW and a $w_{0}=500$ nm TE10 Hermite-Gaussian (HG) mode. For larger $\bar{n}$, the wave function samples a larger field region, suggesting different cooling behaviors when outside/inside the LD regime. (c) Two experimental schemes for carrier-free driving using integrated photonics, based on two counter-propagating beams from the same input to form a passively phase-stable SW, or a TE10 mode delivered from a single grating emitter fed by a TE1 guided mode [22]. The cooling beam’s intensity profiles are sketched on the right. (d) The eight-level structure relevant to ${}^{40}\text{Ca}^{+}$ ($S_{1/2},P_{1/2},D_{3/2}$) with lasers for Doppler cooling. $\Delta_{c},\Delta_{r}$ are detuning between $S_{1/2}$ and $P_{1/2}$, $D_{3/2}$ and $P_{1/2}$ respectively at zero magnetic field.
@@
Key limitations to cooling performance arise from couplings that accompany the desired interactions when employing the usual plane-wave-like running wave (RW) fields. In a frame rotating with laser frequency $\omega$, interaction between a RW laser field and an atom represented as a two-level system with ground and excited states $|g\rangle$ and $|e\rangle$, oscillating in a single motional mode at angular frequency $\omega_{m}$, can be described with a dipole interaction Hamiltonian expanded as:
@@
$\langle e|\hat{V}_{\mathrm{dip}}|g\rangle=\frac{\Omega_{0}}{2}e^{i\bm{k}\cdot\bm{\hat{R}}}=\frac{\Omega_{0}}{2}\left(1+i\eta(\hat{a}+\hat{a}^{\dagger})+...\right),$ (1)
@@
for wave vector $\bm{k}$ with $\left\lvert\bm{k}\right\rvert=k_{0}$ in free space; Rabi frequency $\Omega_{0}$; position operator $\bm{\hat{R}}=\bm{u_{x}}x_{0}(\hat{a}+\hat{a}^{\dagger})$ written in terms of the ground-state wavefunction extent $x_{0}=\sqrt{\frac{\hbar}{2m\omega_{m}}}$, the annihilation and creation operators $\hat{a}$ and $\hat{a}^{\dagger}$, and the unit vector $\bm{u_{x}}$; and Lamb-Dicke (LD) parameter $\eta=\eta_{0}\cos(\theta)$, where $\eta_{0}=k_{0}x_{0}$ and $\theta$ is the angle between $\bm{k}$ and $\bm{u_{x}}$. The terms in Eq. (1) proportional to $\eta\hat{a}$ and $\eta\hat{a}^{\dagger}$ describe coherent coupling to red and blue motional sidebands, respectively [11], driven by first-order field gradients desired for cooling; larger scattering rate on the red as compared to the blue sideband leads to cooling (Fig. 1a) [23, 24]. The leading zeroth order term, due to the field at the ion location, represents carrier coupling that does not affect the motional state. Since typically $\eta<1$, the carrier drive strength exceeds that of the sidebands and is problematic for two reasons. As Rabi frequency is increased to cool more rapidly, this carrier excitation saturates the internal population, reducing contrast in the excitation rate between the red and blue sideband, thereby limiting cooling rates and compromising the final phonon number achievable (Fig. 1a). Additionally, while carrier excitation is associated with no motional excitation, the subsequent spontaneous decay results in net heating. Thus, carrier excitation limits achievable cooling rate and additionally causes net heating.
@@
These limitations can be alleviated via suppression of carrier excitation and selective sideband driving achievable at appropriate locations in structured light fields [25, 26, 27, 28]. For electric dipole transitions of atoms trapped at points with zero intensity and nonzero field gradient, e.g. at SW nodes or nodal lines of higher-order Hermite-Gaussian (HG) modes, the carrier term along with all even orders of the expansion in Eq. 1 is nulled. Fig. 1b gives an example with ${}^{40}\text{Ca}^{+}$ and its 397397397{}397 nm dipole transition. The first-order sideband term then becomes the leading contribution in the Hamiltonian, i.e.
$\langle e|\hat{V}_{\mathrm{dip}}|g\rangle=\frac{\Omega_{0}}{2}\sin(\bm{k}\cdot\bm{\hat{R}})=\frac{\Omega_{0}}{2}\left(\eta(\hat{a}+\hat{a}^{\dagger})+...\right).$ (2)
Note that with this definition the total laser power in the SW configuration is half that of the RW configuration for the same $\Omega_{0}$.
@@
Recent development of integrated waveguide photonics for optical control in planar ion traps [29, 30, 31, 28, 32, 33, 34] offers a route to realize these mechanisms without the phase drifts and pointing instability that pose significant challenges in conventional free-space implementations [35]. For example, SW generation straightforwardly leverages counterpropagating beams derived from a single waveguide input (Fig. 1c, top) as demonstrated in [28]. A TE10 HG mode with transverse field gradients at the intensity node along the beam center (Fig. 1c, bottom) can be achieved via emission of higher-order waveguide modes [22]. While the form of the interaction Hamiltonian is the same to lowest order for these two cases, the effective LD parameter in the latter case, $\eta_{\text{HG}}=\frac{2\sqrt{2}}{w_{0}}x_{0}$ [36] reflects motional coupling from transverse field gradients, additionally offering avenues to engineer interactions to accommodate different ion crystal configurations and motional mode orientations. Note that structured light profiles discussed here can also be used to improve trapped-ion operation beyond laser cooling [37]. While we focus on these relatively simple field configurations, similar considerations would apply to more general structured light profiles [38].
@@
In this work, we examine the practical performance of Doppler and EIT cooling using structured light fields through quantum optical master equation simulations. We consider ${}^{40}\text{Ca}^{+}$’s relevant eight-level structure ($S_{1/2},P_{1/2},D_{3/2}$) in Fig. 1d, although the methods analyzed are applicable to other species. Our treatment extends previous theoretical work in considering Doppler cooling from well beyond the LD regime (Fig. 1b), as well as accounting for sideband saturation effects at high drive strengths for both Doppler and EIT cooling.
@@
In the results presented below we show significant enhancement possible for carrier-free EIT ground-state cooling, simultaneously in cooling rate, motional mode frequency bandwidth, and steady-state phonon number achievable, with over an order of magnitude improvement predicted for the latter. We quantify the sensitivity of these enhancements over typical RW schemes to likely experimental limitations such as polarization and mode-delivery impurities. Realization of these simultaneous advantages may significantly reduce the sideband cooling required to achieve ground-state occupancies desired for high-fidelity optical control. For Doppler cooling, despite the expected degradation in SW cooling of highly excited ions as compared to within the LD regime, we still find that a modest enhancement for SW cooling obtains. While sub-Doppler cooling via polarization gradient cooling can also leverage integrated photonic beam delivery [39] and benefit from ion positioning at polarization SW nodes, the expected enhancement in steady-state phonon number compared to with RWs is limited to 2×2$\times$2 × [40]. Given the more significant gains predicted for EIT here, we focus on EIT as a sub-Doppler and ground-state cooling method. We present our results for the simpler Doppler cooling first followed by EIT cooling, commenting on the impact of $\eta_{0}$ on both schemes. We demonstrate that carrier-free Doppler and EIT cooling use half of the time of RW Doppler and EIT cooling to sequentially cool the ion from outside the LD regime to near ground state with $\bar{n}<10^{-2}$, without further resolved sideband cooling that usually takes multiple milliseconds. This work indicates potential for scalable photonic hardware to alleviate bottlenecks in trapped-ion control, focusing here on laser cooling, and informs experimental implementation in integrated devices.
@@ I Carrier-free Doppler cooling
The treatment of [25] showed that Doppler cooling for ions within the LD regime at SW nodes allows higher cooling rates via use of higher cooling beam intensities, as well as lower steady-state phonon number $\bar{n}_{ss}\equiv\lim_{t\rightarrow\infty}\bar{n}$, than possible with RWs. In a RW configuration and in the LD regime, saturation sets in when $\Omega_{0}\sim\Gamma$. At a SW null, since only sideband couplings are present, saturation instead occurs when $\eta\sqrt{n}\Omega_{0}\sim\Gamma$. Hence, the cooling rate $W_{c}\equiv-\dot{\bar{n}}/\bar{n}\propto(\eta\Omega_{0})^{2}$ can be increased by of order ${\sim 1}/\eta^{2}n$ as compared to with RW fields. This delayed saturation enables a significant gain in $W_{c}$ for sufficiently low phonon numbers $n$ given that typically, $\eta\sim 0.1$. Furthermore, an approximately 2×2$\times$2 × reduction in final phonon number $\bar{n}_{ss}$ is predicted.
@@
However, drastically different behavior, including heating, appears at other positions in the SW profile [25]. Wavefunctions of highly excited ions outside the LD regime, of practical interest to Doppler cooling, span lengthscales comparable to the wavelength (Fig. 1b), and hence, compromised SW cooling performance is expected as compared to within the LD regime as treated in [25]. To study the cooling behavior beyond the LD regime, here we employ numerical Monte-Carlo simulation of the Lindblad master equation and include higher-order expansion terms (see Appendix). Our simulation quantifies the enhancement of Doppler cooling using SWs as compared to RWs both outside and inside the LD regime, including effects of saturation from sideband excitation. We also discuss experimental configurations of carrier-free Doppler cooling based on integrated photonic addressing [28, 31].
@@
In particular, we simulate the eight-level ${}^{40}\text{Ca}^{+}$ structure as an example with a $\hat{\pi}$-polarized cooling beam ($\Omega_{c}$, $\Delta_{c}$) addressing the $S_{1/2}\leftrightarrow P_{1/2}$ transition and a repumper ($\Omega_{r}=2\pi\times 10$ MHz, $\Delta_{r}$) with polarization $\hat{\epsilon}_{p}=\hat{\pi}/\sqrt{2}+(\hat{\sigma}_{+}-\hat{\sigma}_{-})/2$ coupling between $D_{3/2}\leftrightarrow P_{1/2}$ as shown in Fig. 1a, and with a quantizing magnetic field of $B=1$ G. Time evolution of the expectation phonon number $\bar{n}(t)$ starting from a thermal state $\bar{n}(0)$ is extracted from Monte-Carlo simulation of the master equation. The cooling rate at a given $\bar{n}$ is defined as $W_{c}$, and the cooling limit $\bar{n}_{ss}$ is the final steady-state phonon number.
@@
Figure 2: (a) Cooling rate (lines) and limit (points) for $\omega_{m}=2\pi\times 3$ MHz and $\Delta_{c}=-\Gamma/2$. The cooling rate EF is $1.5$ for $\bar{n}=50$ (solid line), and $4.9$ for $\bar{n}=5$ (dotted line). (b) SW and RW cooling trajectory for $\omega_{m}=2\pi\times 3$ MHz. Using $\Delta_{c}=-0.5\Gamma$ and starting at $\bar{n}=50$, SW Doppler cooling reaches $\bar{n}=4$ within $34~{}\mu s$ with Rabi frequency $\Omega_{c}=1.8\Gamma$, while RW Doppler cooling takes $75~{}\mu s$ with $\Omega_{c}=0.8\Gamma$.
@@
Fig. 2a shows $W_{c}$ in the full eight-level system as a function of $\Omega_{c}$ for $\omega_{m}=2\pi\times 3$ MHz. $\Delta_{c}$ is set to $-\Gamma_{c}/2$ which optimizes the cooling limit $\bar{n}_{ss}$. For small $\Omega_{c}$, RW and SW Doppler cooling rates increase quadratically with $\Omega_{c}$ both inside ($\bar{n}=5$) and outside the LD regime ($\bar{n}=50$) before saturation. We define a cooling rate enhancement factor ($\mathrm{EF}$) as a ratio of the maximum achievable cooling rate in either case, i.e. $\mathrm{EF}\equiv W_{SW\max}/W_{RW\max}$. For higher $\bar{n}$, this factor decreases but remains greater than 1, with the maximum cooling rate achieved around the same $\Omega_{c}$ both outside and near the LD regime. Combined with the improvement in $\bar{n}_{ss}$, this validates the applicability and advantage of SW Doppler cooling even for initially highly excited ions. Fig. 2b shows the simulation results of a complete Doppler cooling process from $\bar{n}=50$, outside the LD regime, to the Doppler limit, using the optimal parameters for the cooling rate found in Fig. 2a. We observe that SW Doppler cooling takes less than half of the time compared to RW to reach $\bar{n}=4$, a reasonable starting point for ground-state cooling.